Supposed that a certain ubstance has a half- life of 20 years.If there are presently 2500-milligrams of substance ,then the equation Q(t)=2500(2)−t/20 yields the amount remaining after t-years.How much remains after a. 81 years b.40 years c. 50 years
1
Expert's answer
2019-12-03T11:17:25-0500

a.


"Q(81)=2500\cdot 2^{-\frac{81}{20}}=151\ mg"

b.


"Q(40)=2500\cdot 2^{-\frac{40}{20}}=625\ mg"

c.


"Q(50)=2500\cdot 2^{-\frac{50}{20}}=442\ mg"